OpenStudy (anonymous):

You roll 2 dice. What is the probability that the sum of the dice is odd or 1 die shows a 5? A 6 X 6 table of dice outcomes will help you to answer this question.

OpenStudy (anonymous):

@mary.rojas

OpenStudy (anonymous):

@ybarrap

OpenStudy (anonymous):

@genius12

OpenStudy (mary.rojas):

hope this helps: This is a basic problem: The sum of numbers can be either even or odd.... The probability of having a sum as odd is 1/2. Probability of none of the dices showing a four is (5/6)*(5/6) = 25/36 Thus probability of at least one four is 1 - 25/36 = 11/36 Total probability = 1/2 + 11/36 = 29/36

OpenStudy (anonymous):

yeah that's what we did earlier...

OpenStudy (mary.rojas):

@bahrom7893 help?

OpenStudy (mary.rojas):

so thats not it?

OpenStudy (anonymous):

so I got the pairs: 2,5 5,2 6,5 5,6

OpenStudy (bahrom7893):

1, 2 1, 4 1, 6 2, 1 (disregard) 2, 3 2, 5 3, 2 (disregard) 3, 4 3, 6 4, 1 4, 3 (disregard) 4, 5 5, 2 (disregard) 5, 4 (disregard) 5, 6 6, 1 (disregard) 6, 3 (disregard) 6, 5 (disregard) Those are the ones with an odd sum.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

but not all of them show a 5 like it asks u to

OpenStudy (bahrom7893):

I disregarded the ones that were repeated. And I'm not done yet.

OpenStudy (bahrom7893):

1 die shows a 5: 1,5 2,5 (disregard) 3,5 4,5 (disregard) 5,5 6,5 (disregard)

OpenStudy (bahrom7893):

So we have: 13/36 possible outcomes

OpenStudy (anonymous):

ok now w/out typing "DISREGARD" for some, can u list all the pairs that would go w/ this problem?

OpenStudy (anonymous):

ill start off and u can add any if I miss them

OpenStudy (anonymous):

2,5 5,2 6,5 5,6

OpenStudy (bahrom7893):

1, 2 1, 4 1, 6 2, 3 2, 5 3, 4 3, 6 4, 1 4, 5 5, 6 1, 5 3, 5 5, 5

OpenStudy (anonymous):

oh it says OR huh?

OpenStudy (anonymous):

that why I only listed a few.... my bad

OpenStudy (bahrom7893):

Yea, so you add the possible outcomes.

OpenStudy (anonymous):

ok so that's the whole list??

OpenStudy (bahrom7893):

im pretty sure it is.

OpenStudy (anonymous):

so there's 13...

OpenStudy (anonymous):

so 13/36 ???

OpenStudy (bahrom7893):

There's probably a smarter way of doing this, but i think so.

OpenStudy (bahrom7893):

It's late, and I'm tired after work. Someone should check this over.

OpenStudy (anonymous):

yea thanks bahrom. @oldrin.bataku can u pls check this???

OpenStudy (anonymous):

THE ANSWER WAS WRONG!!!!!!! D: D:

OpenStudy (bahrom7893):

Sorry.. hopefully oldrin will sort things out. I'm falling asleep here.

OpenStudy (bahrom7893):

AND WE HAVE A WINNER!

OpenStudy (anonymous):

The easiest method is just to write out all the possible sums:

OpenStudy (anonymous):